if a collection of cows and crows have 43 heads and 120 feet, how many of each are there?
Answer by checkley71(8403)

COWS+DROWS=43 COWS=43-CROWS
4COWS+2CROWS=120 NOW REPLACE THE COWS BY (43-CROWS)
4(43-CROWS)+2CROWS=120
172-4CROWS+2CROWS=120
-2CROWS=120-172
-2CROWS=-52
CROWS=-52/-2
CROWS=26 ANSWER.
C0WS+26=43
COWS=43-26
COWS=17 ANSWER.
PROOF
4*17+2*26=120
68+52=120
120=120